OpenStudy (anonymous):

Triangle SDF is similar to triangle RTY with ratio of similitude of 4:9 . Also, SF is 2 less than RT and TY is 9 more than RY. IF SD = 8, what is DF?

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I thought we solved it yesterday, right? Not exactly the same, but the same topic

OpenStudy (anonymous):

Yeah we did... But I don't know how to set up the equation..

OpenStudy (loser66):



OpenStudy (loser66):

look at it and let me know if you get why I label them like that. I will guide you more

OpenStudy (anonymous):

Ok.. I understand

OpenStudy (loser66):

The ratio is \(\dfrac{SDF}{RTY}=\dfrac{4}{9}\), that is \(9SDF= 4RTY\) ok?

OpenStudy (anonymous):

Alrighty

OpenStudy (loser66):

Now, for a: we have 9*8 = 4 *a (from the ratio), hence a =??

OpenStudy (loser66):

don't get ???

OpenStudy (anonymous):

Yeah i get it

OpenStudy (anonymous):

now what

OpenStudy (loser66):

a=??

OpenStudy (anonymous):

18

OpenStudy (loser66):

then a-2 =?

OpenStudy (anonymous):

16

OpenStudy (loser66):

ok, now again, a-2 is from the small triangle, and we know that the ratio is 4/9 , that is 9 small = 4 big the associate side of a-2 (from small one) is b (from big one) hence again, 9small = 4 big that is 9(a-2) = 4b, we know that a-2 =16, hence b =???

OpenStudy (loser66):